Shown here is an income statement in the traditional format for a firm with a sales volume of 15 000 units format for a Revenues Cost of goods sold (39,000 $2 95/unit) Gross proft $150,000 $% 760 Seling ($2,150 $0 90/unit) Administration ($5,200+$0.40unit) 15,650 Operating income S 69,900 Required: a. Prepare an income statement i the contribution margin format Variable expenses Total vaniable expenses Fixed expenses Total fixed expenses Refer to your answer to part a when total revenues were $150,000. d-1. Cabaiate the finns operating roome ผู้0r loss)if unit seling prioe and vanate expense per unit do not change and total se onues increase by $14.500. (Do not round intermed ate d-2, Cablate the 5tms operating roome 4or loss)if unt seling poe and varatie expense per unit do not change and total tevenues decrease by $8.00 (Do not round antermediate
Answer #1
a
Contribution margin Income Statement
Revenues 150000
Variable expenses:
Cost of goods sold 44250 =15000*2.95
Selling expenses 13500 =15000*0.9
Administrative expenses 6000 =15000*0.4
Total variable expenses 63750
Contribution margin 86250
Fixed expenses:
Cost of goods sold 9000
Selling expenses 2150
Administrative expenses 5200
Total Fixed expenses 16350
Operating income 69900
b
Contribution margin per unit 5.75 =86250/15000
Contribution margin ratio 57.50% =86250/150000
c-1
Contribution margin 115000 =20000*5.75
Less: Fixed expenses 16350
Operating income 98650
c-2
Contribution margin 40250 =7000*5.75
Less: Fixed expenses 16350
Operating income 23900
d-1
Contribution margin 94587.50 =(150000+14500)*57.5%
Less: Fixed expenses 16350
Operating income 78237.50 or 78238
d-2
Contribution margin 81650 =(150000-8000)*57.5%
Less: Fixed expenses 16350
Operating income 65300